OpenStudy (jennychan12):

Find lim x--> infinity of (3/n)((3i+3)/n)^2

OpenStudy (jennychan12):

\[\lim_{x \rightarrow \infty} \sum_{i = 1}^{n}(\frac{ 3 }{ n })(\frac{ 3i+3 }{ n })^2\]

OpenStudy (jennychan12):

sorry forgot the summation in the question

OpenStudy (anonymous):

Do you mean, as n goes to infinity?

OpenStudy (anonymous):

There is no 'x' in the expression?

OpenStudy (jennychan12):

oh yeah sorry

OpenStudy (anonymous):

I would start by expanding the expression. You can pull out any constants, as well as \(n\) from the summations.

OpenStudy (anonymous):

Then recall that \[ \sum_{i=1}^n i = n(n+1)/2 \\ \sum_{i=1}i^2=n(n+1)(2n+1)/6 \]

OpenStudy (anonymous):

@jennychan12 get it?

OpenStudy (jennychan12):

i'm working on it...

OpenStudy (jennychan12):

i'm expanding the inside for now. i didn't write lim n--> infinity or summation (i = 1 to n) \[\frac{ 3 }{ n }(\frac{ 3i+3 }{ n })^2 = \frac{ 3 }{ n }(\frac{ 9i+18i+9 }{ n^2 })\] \[\frac{ 27i+54i+9 }{ n^3 } = 27(\frac{ (\frac{ n(n+1) }{ 2 }) }{ n^3 }) + 54(\frac{ \frac{ n(n+1)(2n+1) }{ 6 } }{ n^3 })+9(\frac{ n }{ n^3 })\] correct? or did i do something wrong?

OpenStudy (jennychan12):

whoops the last term should be 27(...)

OpenStudy (anonymous):

If you type \left( \right) it will size the parenthesis nicely usually.

OpenStudy (anonymous):

Now you just need to simplify the expression a bit more, so you can take the limit.

OpenStudy (jennychan12):

i did and i got the answer to be 54

OpenStudy (anonymous):

The problem is that I don't understand because you should have \(i^2\) and I don't see it anywhere.

OpenStudy (jennychan12):

oh sorry that should be the first term. oops. i plugged it in wrong. but i did it right on my paper. and i got 54..

OpenStudy (jennychan12):

but the answer is 9 so i did something wrong...

OpenStudy (anonymous):

Try again. I can tell you where you went wrong if I can see your steps.

OpenStudy (jennychan12):

yeah i got the right answer. i forgot to divide 54 by 6 :/